An investor invested a total of $2300 in two mutual funds. One fund earned a 6% profit, and the other a 2% profit. If the total profit was $98, how much was invested in each mutual fund.

Answer by checkley79(3341)

.06X+.02(2,300-X)=98
.06X+46-.02X=98
.04X=98-46
.04X=52
X=52/.04
X=$1,300 WAS INCVESTED @ 6%
2300-1300=$1,000 INVESTED @ 2%.
PROOF:
.06*1300+.02*1000=98
78+20=98
98=98